# The Illusion of State in State-Space Models

### v1

@@ 3 State Tracking
Informally, a state-tracking problem is a problem where the text specifies some sequence of updates to the state of the world, and the goal of the problem is to say what the resulting world state is after the updates have been applied in sequence. This circuit complexity view on the power of neural networks can be combined with other insights from algebraic formal language theory to analyze the kinds of state tracking that SSMs can express. In particular, this theory comprehensively shows us which kinds of state-tracking problems are (likely) not in $\mathsf{TC}^{0}$. This will, in turn, allow us to find examples of hard state tracking that models like SSMs will not be able to solve.
@@ 3.1 State Tracking as a Monoid Word Problem
From the perspective of algebraic formal language theory, state tracking over a finite world can be captured as a word problem on a finite monoid (Liu et al., 2023). [^4] We consider only finite monoids for simplicity, but, in principle, it would be possible to extend this approach to infinite (e.g., finitely generated) monoids as well. Different updates to the world become different elements in the monoid, and resolving the final world state after all the updates have been applied is equivalent to computing the product of a sequence of elements (also called a “word”).
@@
Consider the monoid $\{0,1\}$ where $\cdot$ is addition modulo 2. The word problem involves computing the parity of a string, e.g., $0011\mapsto 0$. From a state-tracking perspective, this monoid captures a world with a single light switch. The identity 00 corresponds to no action whereas $1$ is an update that flips the switch.
Modeling state tracking with word problems lets us draw connections between circuit complexity and abstract algebra to understand which word problems are “hard” to solve. Krohn & Rhodes (1965) established that not all word problems are created equal: some, like Example 3.2, are in $\mathsf{TC}^{0}$, while others are $\mathsf{NC}^{1}$-complete, requiring recurrent processing to solve (Immerman & Landau, 1989; Barrington, 1989). Because we will show SSMs can be simulated in $\mathsf{TC}^{0}$, it follows that $\mathsf{NC}^{1}$-complete state-tracking problems cannot be solved by SSMs (cf. Figure 2).
Whether or not a word problem is $\mathsf{NC}^{1}$-complete depends on the algebraic structure of the underlying monoid. [^5] We focus on word problems on groups, which are monoids with inverses. Barrington (1989) showed that the word problem of every finite non-solvable [^6] Formally, a group $G$ is solvable exactly when there is a series of subgroups $1=G_{0}<G_{1}<\cdots<G_{k}=G$ such that $G_{i-1}$ is normal in $G_{i}$ and $G_{i}/G_{i-1}$ is abelian. group is $\mathsf{NC}^{1}$-complete. That non-solvable groups have $\mathsf{NC}^{1}$-complete word problems is notable because of the ubiquity with which non-solvable groups show up in tasks involving state tracking. The canonical example of an $\mathsf{NC}^{1}$-complete word problem is that of $S_{5}$, the symmetric group on five elements that encodes the permutations over five objects. As an immediate instantiation of this, consider a document describing arbitrarily many sequences of transpositions: “swap ball 1 and 3, swap ball 3 and 5, swap ball 4 and 2, …” [^7] Without less of generality, any permutation can be factored into a sequence of transpositions, or swaps. This means the transpositions over five elements are a generator for $S_{5}$.. Being able to answer the question “where does ball 5 end up?” requires solving the $S_{5}$ word problem. Beyond permutations, Figure 1 shows how many natural state-tracking problems like tracking chess moves, evaluating code, or tracking entities also encode the structure of $S_{5}$, meaning these state-tracking problems also cannot be expressed by a model in $\mathsf{TC}^{0}$. Rather, in order to solve these problems, the depth of the model would have to be expanded to accommodate longer inputs.
@@ 3.2 Encoding $S_{5}$ in Chess State Tracking
Figure 1 already gives some intuition into how state-tracking problems encode $S_{5}$. Out of these examples, the most intricated case is chess. We now give a proper reduction from $S_{5}$ to tracking chess moves, showing formally that not just $S_{5}$, but chess state tracking as well, is $\mathsf{NC}^{1}$-complete.
We define the chess state-tracking problem as follows:
  • Input: The state of a chessboard as well as a sequence of chess moves, where each move is specified as a tuple (source square, target square). Note that this differs from the standard notation that represents a move as a piece along with target square and potential disambiguating information.
  • Output: The resulting board state after starting in the initial board state and applying the sequence of moves one after another, ignoring draws. If any move is illegal given its intermediate board state, we enter a special null board state.
We show that $S_{5}$ can be reduced to chess state tracking, establishing the $\mathsf{NC}^{1}$-completeness of chess state tracking. In other words, we can map any $S_{5}$ sequence to a sequence of chess moves and read off the answer to the $S_{5}$ instance from the final chessboard state.
@@
$S_{5}$ can be reduced to chess state tracking via $\mathsf{NC}^{0}$ reductions.
@@
The idea, as illustrated in Figure 1, is to map each $S_{5}$ element to a sequence of chess moves that permutes five pieces on the chessboard. Then, the final chessboard state will allow us to determine the composition of the permutation sequence. We defer a detailed proof to Appendix B. ∎
Since $S_{5}$ is $\mathsf{NC}^{1}$-complete under $\mathsf{AC}^{0}$ reductions:
@@
Similar reductions can be constructed for evaluating Python or tracking entities in a dialog, as suggested in Figure 1. For another example of a similar reduction used to prove $\mathsf{NC}^{1}$-completeness, we refer the reader to Theorem 3.2 of Feng et al. (2023).
@@ 4 SSMs Can be Simulated in $\mathsf{TC}^{0}$
In this section, we show that the convolutional form of an SSM can be simulated in $\mathsf{TC}^{0}$. Assuming the convolutional form of the model computes the same function as the recurrent form, this implies that SSMs, in whatever parameterization, cannot solve inherently sequential problems, despite their appearance of recurrence and statefulness. We first show containment in $\mathsf{TC}^{0}$ for the simple non-gated variant of generalized S4 models (Theorem 4.1), and then show that the proof goes through for generalized S6 models as well (Theorem 4.2). The main idea in both proofs is that matrix powering can be computed in $\mathsf{TC}^{0}$ (Mereghetti & Palano, 2000), and computing the convolutional form of an SSM can essentially be reduced to matrix powering.
4.1 Simulating Linear SSMs with a $\mathsf{TC}^{0}$ Circuit Family
Theorem 4.1.
For any log-precision SSM $M$ with the S4 architecture, there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes the same function as $M$’s convolutional form.
@@ 4.1 Simulating Linear SSMs with a $\mathsf{TC}^{0}$ Circuit Family
It suffices to construct an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family to simulate a single layer. Then, the full SSM can be simulate by generating this circuit multiple times, routing the output bits from one layer as the inputs to the next. This can be done with log-space overhead by simply storing a counter that tracks the index of the current input gates.
Recall that the S4 convolutional form Eqn. (2) is $h_{i}=\sum_{j=1}^{i}A^{i-j}Bx_{j}$. Crucially, we use the fact that matrix powering over floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Lemma 2.4, extending Mereghetti & Palano, 2000) to print, for each $i,j$, a $\mathsf{TC}^{0}$ circuit family that computes $\pi_{ij}\triangleq A^{i-j}$. Next, we use the fact that fixed-arity arithmetic over floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ to print, for each $i,j$, a circuit that computes
$\tau_{ij}\,\triangleq\,\pi_{ij}Bx_{j}\,=\,A^{i-j}Bx_{j}.$
Since iterated addition over floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023a, extending Hesse, 2001; Chiu et al., 2001 for integers), we can compute $h_{i}\triangleq\sum_{j=1}^{i}\tau_{ij}$. Finally, we compute $y_{i}$ from $x_{i}$ and $h_{i}$ in $\mathsf{TC}^{0}$ via simple arithmetic operations, completing the proof. ∎
4.2 Simulating S6 with a $\mathsf{TC}^{0}$ Circuit Family
Theorem 4.2.
For any log-precision SSM $M$ with the S6 architecture, there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes the same function as $M$’s convolutional form.
@@ 4.2 Simulating S6 with a $\mathsf{TC}^{0}$ Circuit Family
The convolutional form of the S6 layer is given in Eqn. (4). For each $i$ we print an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit computing $\delta_{i},B_{i}$ and $C_{i}$. Since $A$ is diagonal, iterated matrix multiplication is reducible to iterated scalar multiplication, which is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Lemma 2.4). so we can compute the following in $\mathsf{TC}^{0}$:
$\pi_{ij}\triangleq\prod_{k=j+1}^{i}\exp(\delta_{k}A).$
We then conclude analogously to Theorem 4.1, first computing $\tau_{ij}$ from $\pi_{ij}$ and $B_{i}$, then computing $h_{i}=\sum_{j=1}^{i}\tau_{ij}$, and finally computing $y_{i}$ from $x_{i},h_{i},$ and $C_{i}$ in $\mathsf{TC}^{0}$. ∎

### v2

@@ 3 State Tracking
Informally, a state-tracking problem is a problem where the text specifies some sequence of updates to the state of the world, and the goal of the problem is to determine what the world state is after the updates have been applied in sequence. The circuit complexity view on the power of neural networks can be combined with other insights from algebraic formal language theory to analyze the kinds of state tracking that SSMs can express. In particular, this theory reveals which kinds of state-tracking problems are (likely) not in $\mathsf{TC}^{0}$. This will, in turn, allow us to find examples of hard state tracking that models like SSMs cannot express.
@@ 3.1 State Tracking as a Monoid Word Problem
From the perspective of algebraic formal language theory, state tracking over a finite world can be captured as a word problem on a finite monoid (Liu et al., 2023). [^5] We consider finite monoids for simplicity, but the approach may be extendable to infinite (e.g., finitely generated) monoids. Different updates to the world become different elements in the monoid, and resolving the final world state after all the updates have been applied is equivalent to computing the product of a sequence of elements (also called a “word”).
@@
Consider the monoid $\{0,1\}$ where $\cdot$ is addition modulo 2. The word problem is to compute the parity of a string, e.g., $0011\mapsto 0$. From a state-tracking perspective, this monoid captures a world with a single light switch. Identity 00 corresponds to no action, and $1$ flips the switch.
Modeling state tracking with word problems lets us draw connections between circuit complexity and algebra to understand which word problems are hard to solve. Krohn & Rhodes (1965) established that not all word problems are created equal: some, like Example 3.2, are in $\mathsf{TC}^{0}$, while others are $\mathsf{NC}^{1}$-complete, requiring recurrent processing to solve (Immerman & Landau, 1989; Barrington, 1989). Because we will show SSMs can be simulated in $\mathsf{TC}^{0}$, it follows that $\mathsf{NC}^{1}$-complete state-tracking problems cannot be expressed by SSMs (cf. Figure 2).
Whether or not a word problem is $\mathsf{NC}^{1}$-complete depends on the algebraic structure of the underlying monoid. Barrington (1989) showed that the word problem of every finite non-solvable [^6] We focus on word problems on groups, which are monoids with inverses. Formally, a group $G$ is solvable exactly when there is a series of subgroups $1=G_{0}<G_{1}<\cdots<G_{k}=G$ such that $G_{i-1}$ is normal in $G_{i}$ and $G_{i}/G_{i-1}$ is abelian. group is $\mathsf{NC}^{1}$-complete. That non-solvable groups have $\mathsf{NC}^{1}$-complete word problems is notable because of the ubiquity with which non-solvable groups show up in tasks involving state tracking. The canonical example of an $\mathsf{NC}^{1}$-complete word problem is that of $S_{5}$, the symmetric group on five elements that encodes the permutations over five objects. As an immediate instantiation of this, consider a document describing a sequence of transpositions: “swap ball 1 and 3, swap ball 3 and 5, swap ball 4 and 2, …”. Being able to answer the question “where does ball 5 end up?” for all possible swap sequences requires solving the $S_{5}$ word problem. [^7] W.l.o.g., any permutation can be factored into a sequence of transpositions, or swaps. Beyond permutations, Figure 1 shows how many natural state-tracking problems like tracking chess moves, evaluating code, or tracking entities also encode the structure of $S_{5}$, meaning these state-tracking problems also cannot be expressed by a model in $\mathsf{TC}^{0}$. Rather, in order to solve these problems, the depth of the model would have to be expanded to accommodate longer inputs.
@@ 3.2 Encoding $S_{5}$ in Chess State Tracking
Figure 1 already gives some intuition into how state-tracking problems encode $S_{5}$. Out of these examples, the most intricated case is chess. We now give a proper reduction from $S_{5}$ to tracking chess moves, showing formally that not just $S_{5}$, but chess state tracking as well, is $\mathsf{NC}^{1}$-complete. We define the chess state-tracking problem as follows:
  • Input: A chessboard state and sequence of chess moves, where each move is written in UCI notation as a tuple (source square, target square). This differs from the standard SAN notation that represents other information like piece type (Toshniwal et al., 2021).
  • Output: The resulting board state after starting in the initial board state and applying the sequence of moves one after another, ignoring draws. If any move is illegal given the previous board state, a null state is returned.
We show that $S_{5}$ can be reduced to chess state tracking, establishing its $\mathsf{NC}^{1}$-completeness:
@@
$S_{5}$ can be reduced to chess state tracking in UCI notation via $\mathsf{NC}^{0}$ reductions.
@@
Without loss of generality, we consider the variant of $S_{5}$ where the output is true if and only if the original first element returns to the first position after the given sequence of permutations has been applied.
The idea, as illustrated in Figure 1, is to map each element of $S_{5}$ to a fixed sequence of chess moves that permutes five pieces accordingly on the chessboard. Given an instance of the $S_{5}$ word problem, we will construct an initial board state and a sequence of moves such that the final chessboard state encodes the output of that $S_{5}$ problem instance.
Let $M$ denote the set of chess moves in the UCI, i.e., (source square, target square), notation.
Initial Board State. We construct a chessboard similar to Figure 1 but with a black rook at a8 and black queens at b8 to e8.
Chess Move Sequence. We then construct a finite function $f:S_{5}\to M^{*}$ that encodes a permutation $\pi$ as a sequence of chess moves. We first factor each permutation $\pi$ to a sequence of transpositions $\tau_{1}(\pi)\cdots\tau_{m_{\pi}}(\pi)$. Each transposition $\tau\in T$ can in turn be expressed as a sequence of chess moves analogously to Figure 1. For example, transposing items 1 and 3 can be expressed as the move sequence: (a8, a7), (a1, b1), (c8, c6), (b1, a1), (a7, c7), (a1, b1), (c6, a6), (b1, a1), (c7, c8), (a1, b1), (a6, a8), (b1, a1), which has the crucial property that it transposes a8 with c8. We denote the mapping from transpositions to chess move sequences as $f:T\to M^{*}$. Putting it all together, we have
f(π)=
[…]
To reduce a sequence of permutations $w\in S_{5}^{*}$, we let
f(w)=
[…]
Putting It All Together. We call our oracle for chess state tracking with the constructed initial board state and $f(w)$ as the sequence of chess moves. By construction, we can then return true if and only if the rook is at a8. The reduction can be implemented in $\mathsf{NC}^{0}$ because it is a simple elementwise mapping of the input tokens, and decoding from the output chessboard is a finite table lookup. ∎
As a fun aside, we note that the chess board constructed in the above proof is reachable in a standard chess game. The chess sequences encoding permutation sequences are all valid in the game of chess, except that they ignore the fact that repeated board states in chess technically lead to a draw.
Since $S_{5}$ is $\mathsf{NC}^{1}$-complete under $\mathsf{AC}^{0}$ reductions and $\mathsf{NC}^{0}\subseteq\mathsf{AC}^{0}$, we have:
@@
Theorem 3.2 of Feng et al. (2023) uses a similar reduction to prove formula evaluation is $\mathsf{NC}^{1}$-complete. Reductions can be constructed for evaluating Python or tracking entities in a dialog, as suggested by Figure 1. As for chess, the task formatting for entity tracking affects its hardness. For instance, the formatting used by Kim & Schuster (2023) in their Figure 1 is not $\mathsf{NC}^{1}$-complete, whereas the variant shown in our Figure 1 is. This underscores the value of theory for constructing examples of hard state tracking.
@@ 4 SSMs Can be Simulated in $\mathsf{TC}^{0}$
In this section, we show that the convolutional form of common variants of SSM can be simulated in $\mathsf{TC}^{0}$. Assuming the convolutional form of the model computes the same function as the recurrent form, this implies such SSMs cannot solve inherently sequential problems, despite their appearance of recurrence and statefulness. We first show containment in $\mathsf{TC}^{0}$ for non-gated SSMs (Theorem 4.2), and then show the same holds for diagonal SSMs (Theorem 4.4).
4.1 Conditions for Linear SSMs in $\mathsf{TC}^{0}$
Before characterizing specific SSM architectures, we first show that the complexity of computing transition matrix products essentially determines the complexity of simulating an SSM with a circuit family.
Lemma 4.1.
Let $M$ be a log-precision generalized linear SSM. Then there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes $M$’s convolutional form if:
  1. 1.
    For any integer interval $[j,k]$, the matrix product $\prod_{i=j}^{k}\bar{\mathbf{A}}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ as a function of $\bar{\mathbf{A}}_{j},\ldots,\bar{\mathbf{A}}_{k}$ (to $c\log n$ precision for any $c>0$).
  2. 2.
    For $1\leq i\leq n$, $\bar{\mathbf{A}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i},$ and $\mathbf{D}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ as a function of $\mathbf{x}_{i}$.
@@ 4.1 Conditions for Linear SSMs in $\mathsf{TC}^{0}$
Following the proof structure of Merrill & Sabharwal (2023a), we describe how to construct a log-space bounded Turing machine $T_{M}$ that, given $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as input, prints a circuit that simulates $M$ on this input. We first note that for all processing done before or after an SSM layer (projection, non-linearity, layer norm, etc.), $T_{M}$ can follow known simulations of such operations for transformers (Merrill & Sabharwal, 2023a, 2024) to output a $\mathsf{TC}^{0}$ circuit simulating this processing. We thus focus on simulating an individual SSM layer.
Recall from Definition 2.1 that $M$’s convolutional form requires computing $\mathbf{h}_{i}=\sum_{j=1}^{i}\left(\prod_{k=j+1}^{i}\bar{\mathbf{A}}_{k}\right)\bar{\mathbf{B}}_{j}\mathbf{x}_{j}$ and $\mathbf{y}_{i}=\mathbf{C}_{i}\mathbf{h}_{i}+\mathbf{D}_{i}\mathbf{x}_{i}$. By the second precondition, $T_{M}$ can print a $\mathsf{TC}^{0}$ circuit that computes all matrices involved here. Further, by the first precondition, $T_{M}$ can also print a $\mathsf{TC}^{0}$ circuit that computes the innermost product in the computation of each hidden state $\mathbf{h}_{i}$, namely $\prod_{k=j+1}^{i}\bar{\mathbf{A}}_{k}$. It can now print a $\mathsf{TC}^{0}$ circuit to multiply the resulting product [^8] Let $c\log n$ be the SSM’s precision. We compute $\prod_{k}\bar{\mathbf{A}}_{k}$ to $c^{\prime}\log n$ precision for a large enough $c^{\prime}$ (similar to the proof of Lemma A.5) such that the full product $\left(\prod_{k}\bar{\mathbf{A}}_{k}\right)\bar{\mathbf{B}}_{j}\mathbf{x}_{j}$ is correct to at least $c\log n$ bits, as technically required by Definition A.3. with $\bar{\mathbf{B}}_{j}$ and $\mathbf{x}_{j}$, and then print a circuit to compute an iterated sum over the $i$ resulting vectors to compute $\mathbf{h}_{i}$ (cf. iterated addition in Appendix A). It can similarly print a (simpler) circuit to compute $\mathbf{y}_{i}$. Thus, the entire SSM layer can be simulated by an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit. ∎
We will use Lemma 4.1 to show that any non-gated or diagonal generalized linear SSM can be simulated in $\mathsf{TC}^{0}$.
4.2 Non-Gated SSMs are in $\mathsf{TC}^{0}$
Theorem 4.2 (Non-gated SSM).
Let $M$ be a log-precision generalized linear SSM such that, for any $i$,
$\bar{\mathbf{A}}_{i}=\bar{\mathbf{A}},\quad\bar{\mathbf{B}}_{i}=\bar{\mathbf{B}},\quad\mathbf{C}_{i}=\mathbf{C},\quad\mathbf{D}_{i}=\mathbf{D}.$
Then there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes $M$’s convolutional form.
@@ 4.2 Non-Gated SSMs are in $\mathsf{TC}^{0}$
We prove this by showing that both conditions from Lemma 4.1 are satisfied. Computing the matrix product reduces to powering $\bar{\mathbf{A}}^{k-j}$. Crucially, we can use the fact that matrix powering over floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Lemma A.8, extending Mereghetti & Palano, 2000). Finally, $\bar{\mathbf{A}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i},$ and $\mathbf{D}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ because they are constants. ∎
As S4 satisfies the premises of Theorem 4.2, we obtain:
Corollary 4.3.
There exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes S4’s convolutional form.
4.3 Diagonal SSMs are in $\mathsf{TC}^{0}$
Theorem 4.4 (Diagonal SSM).
Let $M$ be a log-precision generalized linear SSM where for $1\leq i\leq n$:
  1. 1.
    the transition matrix $\bar{\mathbf{A}}_{i}$ is diagonal, denoted $\operatorname{diag}(\bar{\mathbf{a}}_{i})$ where $\bar{\mathbf{a}}_{i}\in\mathbb{R}^{d}$;
  2. 2.
    each of $\bar{\mathbf{a}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i}$ and $\mathbf{D}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ as a function of $\mathbf{x}_{i}$.
Then there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes $M$’s convolutional form.
Proof.
By the first condition, $\prod_{i}\bar{\mathbf{A}}_{i}=\prod_{i}\operatorname{diag}(\bar{\mathbf{a}}_{i})$. Iterated multiplication of diagonal matrices is reducible to several iterated scalar multiplications, placing this product in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Lemma A.5). The second condition from Lemma 4.1 is satisfied by assumption. Thus, $M$’s convolutional form is computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
Since S6 satisfies the premises of Theorem 4.4, we have:
Corollary 4.5.
There exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes S6’s convolutional form (used by Mamba).
Proof.
For the first condition, note that S6’s transition matrix $\bar{\mathbf{A}}_{i}$ is defined as $\exp(\delta_{i}\mathbf{A})$ for a fixed diagonal $\mathbf{A}$. The set of diagonal matrices is closed under scalar multiplication and matrix exponentiation, so $\bar{\mathbf{A}}_{i}$ is also diagonal. See Appendix B for a proof that the second condition is satisfied by the S6 parameterization. ∎
Appendix C extends Theorem 4.4 to hold even when $\{\bar{\mathbf{A}}_{i}\}$ are simultaneously diagonalizable, rather than just diagonal. Specifically, we prove the following generalization:
Theorem 4.6 (Simultaneously diagonalizable SSM).
Let $\mathbf{W}$ be a fixed matrix. Let $M$ be a log-precision generalized linear SSM such that, for $1\leq i\leq n$,
  1. 1.
    the transition matrix $\bar{\mathbf{A}}_{i}$ is computable to log precision by the expression $\mathbf{W}\operatorname{diag}(\bar{\mathbf{a}}_{i})\mathbf{W}^{-1}$, where $\bar{\mathbf{a}}_{i}\in\mathbb{R}^{d}$;
  2. 2.
    each of $\bar{\mathbf{a}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i}$ and $\mathbf{D}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ as a function of $\mathbf{x}_{i}$.
Then there exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes $M$’s convolutional form.
This, in turn, allows us to prove that a simultaneously diagonalizable transition matrix generalization of the S6 layer is also in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Corollary C.7).